three years ago, the sum of Mark and Tony's ages was 33. Currently, the sum of Tony and Sam's ages is 36. Sam is 4v years older than Tony. Determine the current ages of Mark's, Tony's, and Sam's.
Answer by ikleyn(52787)

.

            I will solve the problem in two steps.
            Watch my steps.

(1)  From the condition,


        Sam + Tony = 36

        Sam - Tony =  4


    Add the equations. You will get  


        2*Sam = 40,   Sam = 20.


    Hence, Tony is 16 years old.


(2)  Three years ago, the sum of ages Mark + Tony was  33 years.

     Hence, currently it is  33+3+3 = 39 years:


         Mark + Tony = 39,   or

         Mark + 16 = 39.


     Hence, Mark = 39-16 = 23.


ANSWER.  Mark is 23 years old; Tony is 16 years old;  Sam is 20 years old.

Solved.